OpenStudy (anonymous):

Find the common difference for the sequence shown? 1/4, 5/16, 3/8, ...

OpenStudy (anonymous):

5/16 - 1/4 = 5/16 - 4/16 = 1/16 3/8 - 5/16 = 6/16 - 5/16 = 1/16